# Low energy phonons in single crystal ZrW2O8

### v1

@@ Abstract
ZrW2O8 is the prototypical example of a material exhibiting negative thermal expansion (NTE). As well as being among the first NTE materials to be studied in detail, it has served as a test bed for ideas for the physical mechanisms that lead to this unusual behavior. It is now widely accepted that here, and in many other framework materials exhibiting NTE, a collection of low energy phonon modes, as opposed to just one or two, are responsible for the anomalous thermal properties. However, quantitative verification and analysis of the density functional theory (DFT) calculations which underpin this proposal are still lacking. In particular, probing the low energy phonons directly throughout reciprocal space using inelastic neutron scattering, as opposed to other techniques which only probe the Brillouin zone center, is technically challenging and hence rarely done. Here we report inelastic neutron scattering measurements in a large number of Brillouin zones, achieved via the time of flight technique, over a 400 K temperature range. We find excellent agreement between DFT calculations and experimental data at low temperature, albeit with a rescaling of the calculated phonon energies, and of the lattice parameter used in the calculations compared to the measured value. However, the shifts in phonon modes due to the reduction in lattice parameter on warming, that can be predicted using DFT, are not observed. This is most likely due to counteractive anharmonic effects arising from the increased amplitude of lattice vibrations at elevated temperatures. Notwithstanding, the good agreement between experiment and DFT gives high confidence that the theory’s predictions of atomic motions for different modes is correct, and hence can be used as a reliable basis for quantitative analysis of NTE effects.
@@ I Introduction
Negative thermal expansion (NTE) is when a material contracts rather than expands on warming. This effect can be uniform, i.e. along all crystallographic axes [1, 2, 3], or non-uniform [4, 5, 6]. It may also exist over a wide or a more limited range of temperatures [7]. ZrW2O8 is in many respects the canonical example, with its NTE having been re-discovered in the 1990s [8, 9, 10]. It is particularly notable that its NTE behavior is uniform, extant over a wide temperature range, from $\sim 10$ K to $\sim 1000$ K, and its coefficient of thermal expansion is large and negative ($\sim 10^{-5}\rm{K}^{-1}$) (see review articles ref. 11 and ref. 12 and references therein). In common with many other materials known to exhibit NTE, ZrW2O8 has a framework structure, i.e. it contains many ions that form coordinated polyhedra, most of which share corners, with large pores or spaces between the polyhedra. Indeed, it has more recently been found that intercalation of water into these pores can also result in significant anomalous lattice expansion or contraction [13]. The low temperature crystal structure of ZrW2O8 is shown in figure 1.
@@
Figure 1: Crystal structure of ZrW2O8 with ZrO6 octahedra shaded green and WO4 tetrahedra shaded purple and blue, corresponding to tetrahedra centred on the crystallographically distinct W1 and W2 sites respectively. Note that each WO4 tetrahedron contains one unbonded oxygen ion on its vertex, pointing along a local $\langle$1 1 1$\rangle$-direction. The Zr–O–W2 bond angle is $\sim 172^{\circ}$ while the Zr–O–W1 bond angle is $\sim 154^{\circ}$. This figure was created using VESTA [14].
@@
As a material is warmed phonon modes with gradually increasing energy are populated. If the potentials for these modes are asymmetric as a function of displacement, then thermal expansion will arise [15, 16]. Studies of the phonons in NTE materials often involve measurement of the phonon density of states (PDOS) using either Raman spectroscopy, inelastic x-ray scattering, or inelastic neutron scattering measurements of powder samples [17, 18, 19, 20, 21, 22, 23], as this is relatively straightforward to do and can be compared directly to harmonic lattice dynamics calculations such as those from density functional theory (DFT). ZrW2O8 is no exception, with the phonon density of states having been measured as a function of temperature soon after the seminal works on the NTE behavior [24]. It was found that the low energy peaks in the PDOS shift to higher energy on warming. Qualitatively, this is the opposite behavior to normal (positive thermal expansion) in which increased bond lengths on warming result in weaker interatomic potentials and hence a reduction in the energy of phonon modes.
@@
In terms of interpretation of the PDOS measurements, and NTE more generally, early theories concerned the role of rigid unit modes (RUMs), in which Zr–O and W–O polyhedra remain undistorted but the whole structure flexes considerably along their shared vertices. As the amplitude of such modes increases, the root-mean-square distance between metal centers decreases, even though individual metal–oxygen bonds do not change in length [11]. Such modes would typically be found at the lower energy end of the phonon spectrum. However, various measurements have indicated that the picture is somewhat more complex than this. In particular, it has been found from EXAFS [25, 26, 27] that all of the polyhedra in the structure distort somewhat on warming, although the WO4 tetrahedra are noticeably more rigid than the ZrO6 octahedra. Notwithstanding, significant flex occurs along some of the Zr–O–W bonds that comprise the corner-sharing polyhedra. In this context, it would be appealing to search for a particular phonon mode that facilitates this distortion. However, numerous computational studies of the lattice dynamics in ZrW2O8 [28, 29, 30, 31] have demonstrated that there is no such single mode responsible for NTE. Rather, a large number of low energy phonon modes distributed throughout the Brillouin zone have a character that leads to NTE.
@@
The direct measurements of phonons in ZrW2O8 (and many other NTE materials) to date have had a number of limitations. Raman spectroscopy can only measure the $\mathbf{q}=0$ Raman-active modes and thus does not provide a comprehensive picture, although the energy resolution available with this technique is usually superior to that available with neutron or x-ray spectroscopy. Neutron (or x-ray) scattering PDOS measurements are by their very nature a spherical average over reciprocal space of all possible phonon modes at a particular energy, so there is no direct sensitivity to the Q-space origin of shifts in energy with temperature. This can be ameliorated somewhat by combining and interpreting these measurements with DFT calculations. However, this approach is challenging in systems containing many atoms, and hence many phonon modes which may be more closely spaced than the energy resolution of the neutron spectrometer used for the measurements. In contrast, a measurement of a single crystal sample would allow individual modes to be isolated via measurements performed in many different Brillouin zones [32]. This would then provide a more stringent test of the DFT.
@@
The motivation of the present study is therefore to develop a more comprehensive understanding of the low energy phonons in ZrW2O8 , by measuring in more Brillouin zones and with good energy resolution, to identify which particular individual or groups of phonon modes that are especially sensitive to temperature. Such measurements are sensitive to both the eigenvalues and eigenvectors extracted from DFT calculations, and hence pose a stringent test of the range of their validity, most notably for the low energy phonons thought to be important for NTE effects.
@@ II Methods
ZrW2O8 crystallizes in the $P2_{1}3$ cubic space group, with a lattice parameter at room temperature of 9.16 Å[9]. The single-crystalline samples for these experiments were grown as described previously [33]. The sample used was composed of one crystal of mass 2.4 g with a mosaic of $\sim 0.3$ degrees, which is substantially less than the divergence of the incident neutron beam in the inelastic neutron scattering experiments described below, and hence unimportant for the experimental resolution. The lattice parameter was measured to be 9.176 Å  at 2 K, 9.159 Å  at 100 K, 9.1389 Å  at 300 K, and 9.125 Å  at 400 K during the alignment procedure for the inelastic scattering measurements described below. These values are consistent with those reported in the literature for powder samples [9].
II.1 Density functional theory
The phonon dispersion was calculated using Density Functional Perturbation Theory (DFPT) [40], with a $3\times 3\times 3$ Monkhorst-Pack grid of phonon wavevectors and Fourier interpolation of the resulting dynamical matrices, in order to sample the entire Brillouin zone. Additional calculations were performed for fixed unit cells of smaller dimension (9.23, 9.22, 9.215, 9.21 and 9.16 Å), at each of which the structure was relaxed and phonons calculated in the same way as for the fully relaxed structure. The calculated stress for each of these cell dimensions was 0.81, 1.13, 1.29, 1.45 and 3.20  GPa respectively. In the first four of these cases no soft phonon modes were observed, indicating that the computed structure remained dynamically stable. For the calculation with a lattice parameter of 9.16 Å a large number of negative energy modes were observed, indicating that for such a compressed lattice the structure becomes unstable. From this point onwards when we refer to these DFT calculations we mean those with $9.21\leq a\leq 9.23$.
When comparing the computed dispersion with that measured experimentally, an empirical scale factor is applied to the eigenvalues (energies) but not the eigenvectors (structure factors). This is determined by finding the simulation that agrees most closely with the data after such a uniform re-scale: the structure factors from the simulations with different lattice parameters are fitted to the neutron data with a single adjustable parameter (the energy scale). For ZrW2O8 the best overall agreement was found for the case of $a=9.215$and a scale factor $\sim 0.95$. This approach is used because changing the lattice parameter for the simulation does not affect the eigenvalues uniformly.
II.2 Inelastic neutron scattering
The time of flight (ToF) neutron scattering experiments were performed on the IN5 spectrometer at the Institut Laue-Langevin [42]. Data were collected with incident neutron energy $E_{i}=8.24$ meV, with the final chopper spun at 16,000 r.p.m. to give a resolution at the elastic line of 0.27 meV (full-width half-maximum, FWHM). The sample was aligned so that the $(H,H,0)$ and $(0,0,L)$ reciprocal space directions were in the horizontal plane of the instrument. The data were collected by scanning the sample orientation about the vertical axis in 0.5 degree steps over a range of 161 degrees, starting from the $(1,1,0)$-axis perpendicular to the incident neutron beam. Data were collected at temperatures of 100 K and 400 K, with these temperatures chosen in order to increase the signal via the Bose-Einstein population factor while maintaining a large temperature difference between measurements, but avoiding the structural phase transition at $\sim 425$ K [9]. Additional data were collected with the sample aligned with the $(H,0,0)$ and $(0,K,0)$ directions in the horizontal plane of the spectrometer, with the sample orientation scanned over a range of 88 degrees in 1 degree steps. These data were collected with $E_{i}=10.43$ meV and a final chopper speed of 16,000 r.p.m. giving a FWHM energy resolution at the elastic line of 0.35 meV. Data with this configuration were collected at temperatures of 2 K and 295 K.
The raw data were processed (units conversion, normalisation of intensities, etc.) using the Mantid software package [43]. Further processing of the data, and all subsequent visualisation and analysis, including convolution of calculated spectra with the instrumental resolution, were performed using the Horace software package [44]. Calculation of the inelastic neutron scattering cross-section ($S(\mathbf{Q},\omega)$) from the DFT-calculated phonons was done using the Euphonic software package’s Horace plugin [45, 46].
@@ III.1 INS data
Figure 2 shows slices through the 4-dimensional datasets recording $S(\mathbf{Q},\omega)$, in the $\mathbf{Q}$–Energy plane, following a path through reciprocal space traversing a set of high-symmetry points. Note that although the points labelled as X, Y, and Z; and as M1, M2 and M3 are equivalent by symmetry, there are paths between them that are inequivalent by symmetry because the space group, $P2_{1}3$, is non-centrosymmetric. Data are presented centred at $\mathbf{G}=(0,0,4)$ in fig. 2 and $\mathbf{G}=(-2,-2,2)$ in fig. 3 for both $T=100$ K and $T=400$ K, and at $\mathbf{G}=(3,-3,0)$ in fig. 4 for $T=2$ K and $T=295$ K. The DFT-calculated dispersion relation with $a=9.215$, with the energy scale reduced by 5%, is overplotted (see discussion in section II.1 for further details). The calculated structure factor is also shown in panel (c) of all three figures. This set of figures illustrate many of the salient points of this work. It is clear that the dispersion and cross-section calculated using DFT provide an extremely good level of agreement with the neutron measurements, with significant overlap at most wavevectors, indicating that both the eigenvalues and eigenvectors from the DFT are accurate.
@@
Figure 4: Slices through the $\mathbf{Q}$–Energy plane, following the same scheme as fig. 2, with the exception of the extra reciprocal lattice point on the right hand side that is labelled M3 and refers to ($\frac{1}{2},-\frac{1}{2},0)$. The data here were collected in a separate experiment and are centered on the $\mathbf{Q}=(3,-3,0)$ Brillouin zone. Data shown were collected at 2 K and 300 K, which is different to figs. 2 and 3. The different experiments also have a different (but mutually consistent) set of arbitrary units for the intensity scale. These data are notable by having intensity peaked near the G point, unlike the other two datasets shown, and extending to slightly higher energy.
@@
Figure 5: Cut at the mid-point of the Y–R trajectory of fig. 4(a) and (b), i.e. at $\mathbf{Q}=(3.25,-2.5,-0.25)$, showing a significant hardening of the first phonon peak from $\sim 3.5$ meV to $\sim 4$ meV.
@@
Despite the overall excellent agreement between the DFT calculations and the INS measurements shown in figs. 2, 3 and 4, there are a small number of wavevectors for which there is some disagreement. These can be seen by comparing the calculated dispersion and cross-section shown in panel (c) with the data in panels (a) and (b) of all three figures. For example, we can see that for the lowest energy mode on the path M1 – Y (and M2 – Z, which is equivalent by symmetry) the calculated energy is rather higher than the measurement at 100 K, whereas the other low energy ($<4$ meV) modes along other paths agree rather closely at this temperature.
To allow a more quantitative analysis we show constant $\mathbf{Q}$ cuts along the energy axis at various high symmetry wavevectors, G, X, M and R, at different temperatures in fig. 7. In all cases the integration range along the orthogonal $q$-axes is 0.04 reciprocal lattice units. This finite integration range, necessary when plotting data from time-of-flight INS experiments, means that for modes that are highly dispersive near the wavevector of interest the width of peaks may be broader than the nominal instrumental energy resolution. This is most apparent for the single peak in the cut at R – referring to fig. 3 one can see that the dispersion is very steep until quite close to the R point.
@@
Figure 7: Cuts along the energy axis at four high symmetry positions in reciprocal space, centered on different Brillouin zones in order to maximize the cross-section. Panel (a) shows a cut at G, panel (b) a cut at Z, panel (c) a cut at M2, and panel (d) a cut at R. Data at lower temperatures (2 K in panel (a), 100 K in panels (b)–(d)) are shown in blue whereas data at higher temperatures (300 K in panel (a), 400 K in panels (b)–(d)) are shown in red. For all cuts an estimated constant background has been subtracted, followed by correction for the Bose-Einstein population factor in order to allow easier direct comparison of data taken at different temperatures. Solid lines correspond to fits to Gaussians (the number of which is set by the number expected from the DFT calculations), to indicate the peak position.
NTE in ZrW2O8 is reflected in the phonon dispersion relation via the mode Grüneisen parameters, $\gamma_{k}$. For a phonon mode $k$ with frequency $\omega_{k}$ this is defined as
$\gamma_{k}=-\frac{V}{\omega_{k}}\frac{\partial\omega_{k}}{\partial V}$ (1)
where $V$ is the volume of the unit cell. It was by analysis of the mode Gruüneisen parameters obtained from DFT calculations [28, 29, 30, 31] that it was proposed that NTE in ZrW2O8 is dominated by a collection of low energy phonon modes of varying character. However, the way mode Grüneisen parameters are determined from DFT is rather different to what actually happens when warming and cooling a sample. In the DFT, the lattice dynamics are calculated at zero temperature in the quasi-harmonic approximation for a series of different fixed lattice parameters close to the ground state parameters determined from a minimization of the overall forces and energies. This therefore corresponds to the lattice contracting or expanding under pressure / stress, rather than warming / cooling. In reality, the warming will result in an increased amplitude of lattice vibrations, which is not accounted for in the quasi-harmonic approximation, with the consequence that anharmonic effects may become more significant at higher temperatures. This point has been discussed extensively by Mittal et al [30, 12], who have calculated this effect for a few individual modes at particular wavevectors in ZrW2O8 . Some attempt has been made to quantify the effect more broadly using Raman spectroscopy data [17, 18], which has indicated that as well as having large negative Grüneisen parameters, the lowest energy modes also have the largest anharmonicity.
There are two effects which govern the behavior of a phonon mode energy with temperature. On the one hand, in a conventional material, thermal expansion (on increasing temperature) gives rise to longer bonds between atoms which are weaker and hence give rise to lower energy phonon modes. On the other hand, as temperature increases the amplitudes of normal modes increase and intrinsic anharmonicity comes into play. This would then yield phonon modes that harden slightly with increasing temperature [16]. However, in practice in conventional materials this effect is often much smaller than the softening due to thermal expansion and so is masked. In an NTE material one might imagine that if the lattice expansion effect is opposite to normal, then both of the effects described above would add together and result in phonon mode hardening substantially with increasing temperature. Or even in a RUM scenario, in which the bonds between atoms to not change length and hence strength, the intrinsic anharmonicity would be left over to give mode hardening, albeit perhaps smaller in magnitude. Neutron measurements of the phonon density of states in ZrW2O8 [24] indeed indicate that most of the low energy modes in ZrW2O8 do indeed harden on warming. Examining fig. 7, that shows cuts along energy at several key wavevectors in different Brillouin zones, we can gain greater insight into this.
At the G-point (fig. 7(a)) at 100 K, we observe one optic mode at $5.10\pm 0.03$ meV with a FWHM of $0.41\pm 0.06$ meV. On warming to 400 K the mode is fitted to be centered at $5.12\pm 0.05$ meV, with a FWHM of $0.85\pm 0.1$ meV. Our calculations, and those of other groups, indicate that one might expect to see two modes here (see fig. 4, for example). However they are close in energy and dispersive so may not be possible to resolve and may be broadened by additional instrumental effects beyond the bare energy resolution (which is $\sim 0.23$ meV at this energy). Certainly, the peak is noticeably broadened, most likely due to thermal effects, at 400 K. The behavior of this peak contrasts with the expected softening due to the change in lattice alone of -0.2 meV (for both calculated modes), which Gupta et al [30] predict to be compensated such that the expected shifts are $-0.07$ meV and $-0.03$ meV respectively, which is significantly less than the energy resolution of our measurements  [^1] Note that the predictions of Gupta et al [30] refer to a change in temperature of 0 K to 300 K, i.e. the same difference in temperature as for our measurements. The offset is unlikely to be important, since NTE in ZrW2O8 is linear over this temperature range.. The Raman measurements of Oishi et al include this mode [18], which they find to be centered at 5.16 meV and largely independent of temperature, consistent with our finding.
At the Z-point (fig. 7(b), equivalent to the X- and Y-points) we observe three modes. At 100 K they are centered at energies of $3.50\pm 0.05$ meV, $4.43\pm 0.05$ meV and $5.28\pm 0.04$ meV. All three were found to have the same FWHM (within error) of $0.49\pm 0.05$ meV. At 400 K these peaks are centered at energies of $3.70\pm 0.10$ meV, $4.48\pm 0.09$ meV and $5.44\pm 0.13$ meV respectively. The FWHMs were found to be $0.53\pm 0.11$ meV. Gupta et al presented predictions for the first two of these modes, with a shift due to the change in lattice of -0.14 and -0.16 meV respectively, that is compensated by thermal effects to give expected shifts of 0.22 meV and 0.68 meV. The first of these predictions is consistent with our measurement, whereas the latter is not – the predicted effect of temperature (referred to as explicit anharmonicity) is larger than what is actually measured.
At the M-point (fig. 7(c), actually M2 in our notation) we observe three modes. At 100 K they are centered at energies of $3.77\pm 0.04$ meV, $4.35\pm 0.05$ meV and $5.60\pm 0.03$ meV. All three were found to have the same FWHM (within error) of $0.42\pm 0.04$ meV. At 400 K these peaks are centered at energies of $3.58\pm 0.07$ meV, $4.05\pm 0.06$ meV and $5.62\pm 0.04$ meV respectively. The FWHMs were found to be $0.34\pm 0.06$ meV. For the first two of these modes, Gupta et al predict shifts due to a change in lattice of -0.36 and -0.37 meV respectively, compensated by explicit anharmonicity to give overall predicted shifts of 0.06 and 0.18 meV. In contrast, our fits would seem to indicate that both modes soften slightly. However, inspecting fig. 3 we see that care should be taken here – it is not obvious from the colormap that these two modes soften as much as the cuts and fits seem to suggest. In fact, what appears to happen is the lowest energy mode which disperses weakly along the Z – M2 trajectory with an energy of $\sim 3.5$ meV seems to steepen slightly on warming, with the effect of bringing some of its spectral weight into the integration region of the cut at 400 K shown in fig. 7(c). Therefore, in the absence of a calculation for the shifts in energy at all $\mathbf{Q}$, we cannot quantify with much accuracy how well the measurement and prediction agree at this wavevector. The predictions suggest small shifts at about the level of our experimental resolution, which looks qualitatively reasonable.
At the R-point (fig. 7(d)) we observe one mode. At 100 K it is centered at an energy of $4.45\pm 0.02$ meV, with a FWHM of $0.56\pm 0.03$ meV. At 400 K the peak is centered at an energy of $4.43\pm 0.05$ meV with a FWHM of $0.91\pm 0.11$ meV. Here, Gupta et al predict a shift due to the change in lattice of -0.39 meV, which is complemented by a significant explicit anharmonicity of the opposite sign to those at other wavevectors, to give an overall predicted shift of -0.77 meV. The large FWHM observed at both temperatures is due to the steep dispersion of the modes near the R-point and the consequent inclusion in the cut integration region of spectral weight away from the R-point itself. Notwithstanding this, it is clear from both fig. 7(d) and fig. 3 that the mode near R does not soften as much as predicted.

### v2

@@ Abstract
ZrW2O8 is the prototypical example of a material exhibiting negative thermal expansion (NTE). It is now widely accepted that in ZrW2O8, and in many other framework materials exhibiting NTE, a collection of low energy phonon modes, as opposed to just one or two, are responsible for the anomalous thermal properties. However, quantitative verification and analysis of the density functional theory (DFT) calculations which underpin this proposal are still lacking. In particular, probing the low energy phonons directly throughout reciprocal space using inelastic neutron scattering, as opposed to other techniques which only probe the Brillouin zone center, is technically challenging and hence rarely done. Here we report inelastic neutron scattering measurements in a large number of Brillouin zones over a 400 K temperature range. We find excellent agreement between DFT calculations and experimental data at low temperature. However, the shifts in phonon modes predicted by DFT due to the reduction in lattice parameter (warming) are not observed. This is most likely due to counteractive anharmonic effects, which we verified using finite temperature molecular dynamics (MD) calculations. Notwithstanding, both DFT and MD results are consistent with NTE in ZrW2O8 arising from the tension effect, and by extension this explanation is supported by the neutron scattering results.
@@ I Introduction
Negative thermal expansion (NTE) is when a material contracts rather than expands on warming. This effect can be uniform, i.e. along all crystallographic axes [1, 2, 3], or non-uniform [4, 5, 6]. It may also exist over a wide or a more limited range of temperatures [7]. ZrW2O8 is in many respects the canonical example, with its NTE having been re-discovered in the 1990s [8, 9, 10]. It is particularly notable that its NTE behavior is uniform, extant over a wide temperature range, from $\sim 10$ K to $\sim 1000$ K, and its coefficient of thermal expansion is large and negative ($\sim 10^{-5}\rm{K}^{-1}$) (see review articles ref. 11 and ref. 12 and references therein). In common with many other materials known to exhibit NTE, ZrW2O8 has a framework structure, i.e. it contains many ions that form coordinated polyhedra, most of which share corners, with large pores or spaces between the polyhedra (see fig. 1. It has also been found that intercalation of water into these pores can also result in significant anomalous lattice expansion or contraction [13].
@@
Figure 1: The crystal structure of ZrW2O8 (space group $P2_{1}3$) with ZrO6 octahedra shaded green and WO4 tetrahedra shaded purple and blue, corresponding to tetrahedra centred on the crystallographically distinct W1 and W2 sites respectively. Note that each WO4 tetrahedron contains one unbonded oxygen ion on its vertex, pointing along a local $\langle$1 1 1$\rangle$-direction. This figure was created using VESTA [14].
@@
As a material is warmed, phonon modes with gradually increasing energy are populated. If the potentials for these modes are asymmetric as a function of displacement, then thermal expansion will arise [15, 16]. Studies of the phonons in NTE materials often involve measurement of the phonon density of states (PDOS) using either Raman spectroscopy, inelastic x-ray scattering, or inelastic neutron scattering measurements of powder samples [17, 18, 19, 20, 21, 22, 23], as this is relatively straightforward to do and can be compared directly to quasi-harmonic lattice dynamics calculations such as those from density functional theory (DFT). ZrW2O8 is no exception, with the phonon density of states having been measured as a function of temperature soon after the seminal works on the NTE behavior [24]. It was found that the low energy peaks in the PDOS shift to higher energy on warming. Qualitatively, this is the opposite behavior to normal (positive thermal expansion) in which increased bond lengths on warming typically result in weaker interatomic potentials and hence a reduction in the energy of phonon modes.
@@
In terms of interpretation of the PDOS measurements, and NTE more generally, early theories concerned the role of rigid unit modes (RUMs), in which Zr–O and W–O polyhedra remain undistorted but the whole structure flexes considerably along their shared vertices. As the amplitude of such modes increases, the root-mean-square distance between metal centers decreases, even though individual metal–oxygen bonds do not change in length [11]. Such modes would typically be found at the lower energy end of the phonon spectrum. However, various measurements have indicated that the picture may be somewhat more complex than this. In particular, it has been found from EXAFS [25, 26, 27] that all of the polyhedra in the structure distort somewhat on warming, although the WO4 tetrahedra are noticeably more rigid than the ZrO6 octahedra. Notwithstanding, significant flex occurs along some of the Zr–O–W bonds that comprise the corner-sharing polyhedra. In this context, it would be appealing to search for a particular phonon mode that facilitates this distortion. However, numerous computational studies of the lattice dynamics in ZrW2O8 [28, 29, 30, 31] have demonstrated that there is no such single mode responsible for NTE. Rather, a large number of low energy phonon modes distributed throughout the Brillouin zone have a character that leads to NTE.
@@
The direct measurements of phonons in ZrW2O8 (and many other NTE materials) to date have had a number of limitations. Raman spectroscopy can only measure the Brillouin zone center ($\mathbf{q}=0$) Raman-active modes and thus does not provide a comprehensive picture, although the energy resolution available with this technique is usually superior to that available with neutron or x-ray spectroscopy. Neutron (or x-ray) scattering PDOS measurements are by their very nature a spherical average over reciprocal space of all possible phonon modes at a particular energy, so there is no direct sensitivity to the Q-space origin of shifts in energy with temperature. This can be ameliorated somewhat by interpreting these measurements with DFT calculations. However, this approach is challenging in systems containing many atoms, and hence many phonon modes which may be more closely spaced than the energy resolution of the neutron spectrometer used for the measurements. In contrast, a measurement of a single crystal sample would allow individual modes to be isolated via measurements performed in many different Brillouin zones (see appendices for a more detailed discussion). This would then provide a more stringent test of the DFT.
@@
The motivation of the present study is therefore to develop a more comprehensive understanding of the low energy phonons in ZrW2O8 , by measuring in more Brillouin zones using inelastic neutron scattering with good energy resolution, to identify if any particular individual or groups of phonon modes are especially sensitive to temperature. Such measurements are sensitive to both the eigenvalues and eigenvectors extracted from ab-initio calculations, and hence pose a stringent test of the range of their validity, most notably for the low energy phonons thought to be important for NTE.
@@ II Methods
Time of flight neutron scattering experiments were performed on the IN5 spectrometer at the Institut Laue-Langevin [32]. Data were collected with incident neutron energies $E_{i}=8.24$ meV and $E_{i}=10.43$ meV, at temperatures in the range 2 K to 400 K. Density Functional Theory (DFT) calculations were performed using the plane-wave pseudopotential method as implemented in the CASTEP package [33]. Calculations were performed for a series of different fixed lattice parameters close to the ground state parameters determined from a minimization of the overall forces and energies. Phonon dispersions were calculated using Density Functional Perturbation Theory (DFPT). It is important to note that in DFT the lattice dynamics are calculated at zero temperature in the quasi-harmonic approximation. When comparing the computed dispersion with that measured experimentally with neutrons, an empirical scale factor was applied to the eigenvalues (phonon energies) but not to the eigenvectors (structure factors).
Calculations of the finite-temperature behavior of the phonon dispersion at a few high-symmetry wavevectors were performed using machine learned potentials and molecular dynamics (MD). The potentials were computed using the automated potential development (APD) workflow [34]. These potentials were then used for MD simulations, using the LAMMPS software package [35].
Further details of the inelastic neutron scattering experimental methods can be found in the appendices, as can the detailed schemes used for the density functional theory and molecular dynamics calculations.
@@ III.1 INS data
Figure 4: Slices through the $\mathbf{Q}$–Energy plane, following the same scheme as fig. 2, with the exception of the extra reciprocal lattice point on the right hand side that is labeled M3 and refers to ($\frac{1}{2},-\frac{1}{2},0)$. The data here were collected in a separate experiment and are centered on the $\mathbf{Q}=(3,-3,0)$ Brillouin zone. Data shown were collected at 2 K and 300 K, which is different to figs. 2 and 3. The different experiments also have a different (but mutually consistent) set of arbitrary units for the intensity scale. These data are notable by having intensity peaked near the G point, unlike the other two datasets shown, and extending to slightly higher energy.
Figure 2 shows slices through the 4-dimensional datasets recording $S(\mathbf{Q},\omega)$, in the $\mathbf{Q}$–Energy plane  [^1] By convention energy is denoted by frequency $\omega$ in the expression of the neutron structure factor., following a path through reciprocal space traversing a set of high-symmetry points. Note that although the points labeled as X, Y, and Z; and as M1, M2 and M3 are equivalent by symmetry, there are paths between them that are inequivalent by symmetry because the space group, $P2_{1}3$, is non-centrosymmetric. Data are presented centered at $\mathbf{G}=(0,0,4)$ in fig. 2 and $\mathbf{G}=(-2,-2,2)$ in fig. 3 for both $T=100$ K and $T=400$ K, and at $\mathbf{G}=(3,-3,0)$ in fig. 4 for $T=2$ K and $T=300$ K. The DFT-calculated dispersion relation with $a=9.215$, with the energy scale reduced by 5%, is overplotted. The calculated structure factor is also shown in panel (c) of all three figures. This set of figures illustrate many of the salient points of this work. It is clear that the dispersion and cross-section calculated using DFT provide an extremely good level of agreement with the neutron measurements, with significant overlap at most wavevectors, indicating that both the eigenvalues and eigenvectors from the DFT are accurate.
@@
Figure 5: Cut at the mid-point of the Y–R trajectory of fig. 4(a) and (b), i.e. at $\mathbf{Q}=(3.25,-2.5,-0.25)$, showing hardening of the first phonon peak from $\sim 3.5$ meV to $\sim 4$ meV.
@@
Despite the overall excellent agreement between the DFT calculations and the INS measurements shown in figs. 2, 3 and 4, there is some disagreement. An example can be seen by comparing the calculated dispersion and cross-section shown in panel (c) with the data in panels (a) and (b) of all three figures. For example, we can see that for the lowest energy mode on the path M1 – Y (and M2 – Z, which is equivalent by symmetry) the calculated energy is rather higher than the measurement at 100 K, whereas the other low energy ($<4$ meV) modes along other paths agree rather closely at this temperature.
@@
Figure 7: Cuts along the energy axis at four high symmetry positions in reciprocal space, centered on different Brillouin zones in order to maximize the cross-section. Panel (a) shows a cut at G, panel (b) a cut at Z, panel (c) a cut at M2, and panel (d) a cut at R. Data at lower temperatures (2 K in panel (a), 100 K in panels (b)–(d)) are shown in blue whereas data at higher temperatures (300 K in panel (a), 400 K in panels (b)–(d)) are shown in red. For all cuts an estimated constant background has been subtracted, followed by correction for the Bose-Einstein population factor in order to allow easier comparison of data taken at different temperatures. Solid lines correspond to fits to Gaussians (the number of which is set by the number expected from the DFT calculations), to indicate the peak positions.
To allow a more quantitative analysis we show constant $\mathbf{Q}$ cuts along the energy axis at various high symmetry wavevectors, G, X, M and R, at different temperatures in fig. 7. In all cases the integration range along the orthogonal $q$-axes is 0.04 reciprocal lattice units. This finite integration range, necessary when plotting data from time-of-flight INS experiments, obviously means that for modes that are highly dispersive near the wavevector of interest the width of peaks may be broader than the nominal instrumental energy resolution. This is most apparent for the single peak in the cut at R – referring to fig. 3 one can see that the dispersion is very steep until quite close to the R point.